OpenStudy (anonymous):

a man drives to work at 40mph. He drives home,along the same route, at 60mph. What is his average speed overall.

OpenStudy (anonymous):

Suppose he lived 120 miles from work. It would take him 3 hours to get there and 2 hours back. So he'd travel 240 miles in 5 hours. So what is his speed for this entire trip? (divide 240 by 5).

OpenStudy (anonymous):

thanks that souds reasonable\

OpenStudy (zarkon):

\[\left(\frac{1}{40}+\frac{1}{60}\right)/2=\frac{1}{48}\]